OpenStudy (anonymous):

will award medal for help. how do i start this problem? Find the coordinates of the other endpoint when you are given the midpoint (point M) and one of the endpoints (point P). P = (10, 6) and M = (-4, 8) (-18, 10) (2, 10) (3, 7)

OpenStudy (anonymous):

this is an example : http://www.purplemath.com/learning/viewtopic.php?f=11&t=676

OpenStudy (aum):

Assume the other end point has the coordinates (p, q) The midpoint of (10,6) and (p,q) is ( (10+p)/2, (6+q)/2 ) = (-4, 8) (10+p)/2 = -4 10+p = -8 p = -18 (6+q)/2 = 8 6+q = 16 q = 10 (-18,10)